What are the slope and y-intercept of a line that contains the points (5,-1) and has the same y-intercept as x-3y=6?
Answer by lwsshak3(11628)

What are the slope and y-intercept of a line that contains the points (5,-1) and has the same y-intercept as x-3y=6?
**
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
x-3y=6
-3y=-x+6
divide by -3
y=(x/3)-2
y-intercept=-2
..
For line with same y-intercept and containing point (5,-1)
y=mx-2
Using coordinates of given point (5,-1) to solve for m
-1=m*5-2
5m=1
m=1/5
Equation: y=(x/5)-2